OpenStudy (anonymous):

solve for w 2w^2-2=3w

OpenStudy (anonymous):

sorry i didnt see this earlier. you must set it equal to zero by subtracting the 3w over: 2w^2-3w-2=0 the factors are going to look something like: (2w )(w ) then you need to figure out what needs to be added and subtracted to each (2w+1)(w-2)=0 so: 2w+1=0 AND w-2=0 w=-1/2 w=2 ANSWER: w= -1/2, 2